# Diagnosing and Mitigating Modality Interference in Multimodal Large Language Models

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated impressive capabilities across tasks, yet they often exhibit difficulty in distinguishing task-relevant from irrelevant signals—particularly in tasks like Visual Question Answering (VQA)—which can lead to susceptibility to misleading or spurious inputs. We refer to this broader limitation as the Cross-Modality Competency Problem—the model’s inability to fairly evaluate all modalities. This vulnerability becomes more evident in modality-specific tasks—such as image classification or pure text question answering—where models are expected to rely solely on one modality. In such tasks, spurious information from irrelevant modalities often lead to significant performance degradation. We refer to this failure as Modality Interference, which serves as a concrete and measurable instance of the cross-modality competency problem, and we further design a perturbation-based causal diagnostic experiment to verify and quantify this problem. To mitigate modality interference, we propose a novel framework to finetune MLLMs, including perturbation-based data augmentations with both heuristic perturbations and adversarial perturbations via Projected Gradient Descent (PGD), and a consistency regularization strategy applying on model outputs with original and perturbed inputs. Experiments on multiple benchmark datasets (image-heavy, text-heavy and VQA tasks) and multiple model families with different scales demonstrate significant improvements in robustness and cross-modality competency, indicating our method’s effectiveness in boosting unimodal reasoning ability while enhancing performance on multimodal tasks.
@@ 1 Introduction
Recent studies have attributed this phenomenon to a variety of symptoms arising during the vision-language alignment process, such as catastrophic forgetting imagewikiqa ; mof ; cogvlm ; vila , knowledge conflict vlmconflict ; wang2024mdpo , and spurious correlations more ; spurlens . Catastrophic forgetting has been identified as a key factor in visual degradation imagewikiqa ; mof , where multimodal tuning of MLLM overrides its pretrained visual features. Cross-modal knowledge conflict vlmconflict ; wang2024mdpo impairs pure-text reasoning, as models often produce inconsistent outputs when visual inputs are introduced, reflecting misaligned visual and textual parametric memories. Additionally, studies on spurious correlations more ; spurlens show that MLLMs tend to rely on superficial cross-modal cues rather than task-relevant grounding. While these symptoms shed light on MLLMs’ limitations, most works treat these issues in isolation. For instance, architectural issues such as shallow cross-modal fusion have been widely discussed: lightweight projectors in models like LLaVA llava fail to fully align vision and language representations, resulting in unstable modality reliance mof ; vlmconflict . Others attribute performance bottlenecks to data limitations—insufficient modality-specific supervision leads to impaired visual decoding imagewikiqa and diminished language understanding vila . Inspired by these observations, our insight is to unify these challenges under a broader perspective: the model’s inability to identify and rely on the modality that contributes most relevant information to the task. We argue that the fundamental limitation lies in MLLMs’ lack of cross-modality competency Competency —the ability to fairly evaluate and integrate information across modalities. Current MLLMs lack mechanisms to support this competency during inference, making them vulnerable to misleading cross-modal signals—a failure mode we refer to as Modality Interference.
@@
To systematically diagnose and mitigate modality interference, we introduce a two-stage methodology grounded in causal analysis. First, we design a perturbation-based evaluation experiment inspired by causal intervention principles more to diagnose the extent of modality interference across tasks and model scales. Second, we propose a robust fine-tuning framework to mitigate modality interference. Specifically, in our evaluation analysis, we first focus on modality-heavy settings using a multiple-choice question answering format, where the model selects an answer from a fixed set of options based on both image and text input. We then include image-heavy tasks (e.g., image classification), text-heavy tasks (e.g., pure-text QA), and balanced multimodal tasks (e.g., VQA), allowing us to examine how models behave under different modality-reliance scenarios. To further induce modality interference, we introduce heuristic perturbations: In image-heavy tasks, we perturb the text input by prepending either (i) unrelated scientific facts or (ii) misleading descriptions that falsely associate an incorrect option with the image content. In text-heavy tasks, where the default visual input is random noise, we perturb the visual input with (i) semantically meaningful real images, (ii) full black canvas, or (iii) full white canvas. These perturbations are designed to either introduce spurious cues or reinforce irrelevant modality signals. We evaluate the resulting changes in model predictions to assess the robustness of modality selectivity. While the perturbation-based evaluations offer empirical insights, we further frame our analysis through a causal intervention framework and in which we model modality interference through a causal graph abstraction. Building on this framework, we evaluate a range of pretrained MLLMs across different architectures and scales with results shown in Figure 1. In image-heavy tasks, unrelated textual facts moderately reduce performance, while misleading descriptions cause severe degradation—revealing the model’s vulnerability to spurious textual cues. In text-heavy tasks, canvas inputs have little effect, but unrelated real images significantly hurt performance, indicating improper fusion of irrelevant visual signals into reasoning.
@@
The empirical results from Figure 1 confirm the presence of modality interference and reveal the limitations of current MLLMs in lack of cross-modality competency. To mitigate modality interference, we propose a perturbation-based fine-tuning framework for MLLMs. Specifically, to alleviate data insufficiency, we apply a perturbation-based data augmentations strategy, where we construct a diverse mixture of modality-specific, perturbation-augmented samples and original VQA samples. The perturbations includes both heuristic perturbations (e.g. add unrelated facts in image-heavy prompts) and adversarial training time perturbations, which exposes the model to worst-case alignment disruptions, serving as a strong variant. To further improve robustness, we introduce a consistency regularization strategy(e.g., via Jensen–Shannon divergence), which enforces output stability between original and perturbed samples. In summary, the main contributions of this paper are threefold. First, we introduce the notion of the Cross-Modality Competency Problem to describe how multimodal models may struggle to balance different modalities, and analyze modality interference as one concrete instance of such challenges in MLLMs. Second, we design a perturbation-based causal evaluation experiment that systematically quantifies modality reliance and reveals models’ susceptibility to modality interference. Third, we propose a fine-tuning strategy that combines supervised augmentation with both heuristic and adversarial perturbations and consistency regularization to mitigate modality interference. Extensive experiments across multiple MLLM families llava15 ; instructblip ; qwen25vl and diverse benchmarks miniimagenet ; caltech101 ; openbookqa ; mmlu ; scienceqa ; mmbench ; seedbench demonstrate the superiority of our method.
@@ 3 Causal Analysis on Modality Interference
Competency problems describe scenarios where models rely on spurious correlations between isolated input features and output labels to make predictions, instead of leveraging meaningful interactions among multiple features Competency . We extend this concept to the multimodal setting by treating entire modalities (e.g., image $X_{I}$ or text $X_{T}$) as structured feature sources. We define the Cross-Modality Competency Problem as a class of tasks where no single modality should be sufficient to determine the correct prediction, and models are expected to integrate information across modalities to make robust decisions. A competent model should be able to fairly evaluate and integrate all modalities, identifying which ones carry task-relevant signals while ignoring misleading or irrelevant ones. For instance, in a pure-text question answering task, the model receives both a question and an image, as is standard in MLLM input formats. However, the image is not required to answer the question. If the model relies on spurious visual cues—such as objects or scenes that frequently co-occur with certain answers—it violates the task’s competency condition by grounding predictions in irrelevant modality signals. This manifests as Modality Interference, where the presence of an irrelevant but misleading modality disrupts the model’s reasoning.
@@
To systematically measure cross-modality competency, we propose a perturbation-based evaluation framework. The core idea is to inject controlled noise into the irrelevant modality and assess the model’s robustness to such perturbations. Specifically, for image-heavy tasks, we perturb the text input by: (1)Prepending unrelated scientific facts; (2) Prepending misleading descriptions that falsely link incorrect options to the image content. For text-heavy tasks, we perturb the visual input by: (1) Attaching a real but irrelevant image; (2) Substituting with a full black or full white canvas image. Models with strong modality selectivity should maintain high prediction consistency when irrelevant modality signals are perturbed. We select multiple image-heavy and text-heavy tasks for evaluation. Each task is framed as a multiple-choice classification problem, requiring the model to choose the correct option based on image and text modalities as input, with perturbations applied as described above. Details in Section 5 and Section C.1.
@@
We formalize modality interference through a causal intervention perspective with a causal graph, as shown in Figure 2, where visual inputs ($X_{I}$) and textual inputs ($X_{T}$) are processed into their respective representations ($Z_{I}$, $Z_{T}$) before being fused to produce the final prediction ($A$). To study the model’s reliance on different modalities, we introduce perturbations directly at the input level, serving as causal interventions causalIntervention on $X_{I}$ and $X_{T}$. Under ideal cross-modality competency, the model’s prediction should primarily depend on the task-relevant pathway (e.g., $X_{I}\rightarrow Z_{I}\rightarrow A$ in image-heavy tasks, $X_{T}\rightarrow Z_{T}\rightarrow A$ in text-heavy tasks). Causal interventions at the input level allow us to diagnose whether the model improperly fuses irrelevant signals into its decision process. We use $x_{I}\textquoteright$ ’ to denote an intervention on image $X_{I}$ and use $x_{T}\textquoteright$ ’ as an intervention on text $X_{T}$. Following Pearl’s causal framework causalIntervention ; intermediaryVariable , we quantify the impact of modality perturbations on model predictions by formalizing causal effects in our multimodal setting. Specifically, we define the pre-intervention prediction distribution as $P(A|X_{I},X_{T})$, and the post-intervention prediction distribution after applying a perturbation on $X_{I}$ or $X_{T}$ as $P\textquoteright(A|\text{do}(X_{I}=x_{I}\textquoteright)$ ’ ) or $\text{do}(X_{T}=x_{T}\textquoteright))$ ’ ) ). The do-operation represents an intervention to specific modality, and the causal effect (CE) of an intervention is evaluated via a distance metric $\delta$ comparing $P$ and $P\textquoteright$ ’ as $\text{CE}=\delta(P,P^{\prime})$. We assess the causal effect via prediction changes using $\delta_{\text{cp}}(P,P^{\prime}):=\mathbb{I}\left(a\neq a^{\prime}\right)$ in which $a=\arg\max_{x}P(x)$ is the predicted answer before intervention, $a\textquoteright=\arg\max_{x}P\textquoteright(x)$ is the predicted answer after intervention and $\mathbb{I}(\cdot)$ is the indicator function that outputs $1$ if $a\neq a\textquoteright$ ’ and 00 otherwise. Thus, $\delta_{\text{cp}}$ captures whether the model’s final decision $A$ changes under perturbations to the input modality. In all interventions, a high value of $\delta_{\text{cp}}$ indicates the model’s susceptibility to modality interference, revealing spurious reliance on irrelevant modality.
@@ 4 Methods
To mitigate modality interference and enhance cross-modality competency, we propose a unified perturbation-aware training framework that introduces interventions at both the input level (on $X_{I}$ and $X_{T}$) and the representation level (on $Z_{I}$ and $Z_{T}$) with consistency regularization. Overall pipeline is displayed in  Figure 3.
@@ 4.1 Perturbation-based Data Augmentation
We sample $N_{\text{img}}$ and $N_{\text{text}}$ examples from $\mathcal{D}^{\text{img}}$ and $\mathcal{D}^{\text{text}}$ to construct $\mathcal{B}^{\text{img}}_{\text{orig}}$ and $\mathcal{B}^{\text{text}}_{\text{orig}}$ respectively, and the remaining $N_{\text{vqa}}$ examples are VQA samples from $\mathcal{D}^{\text{vqa}}$ to construct $\mathcal{B}^{\text{vqa}}$ With dynamically generated perturbed variants for each image-heavy and text-heavy sample, the final training batch is:
@@
where $\alpha$ is the step size and $\Pi$ projects the noise back into the $\ell_{\infty}$ ball with radius $\epsilon$, we initialize $\delta$ with Gaussian noise $\mathcal{N}(0,\epsilon^{2})$. After obtaining $\tilde{\mathbf{E}}$, we compute the adversarial loss as $\mathcal{L}_{\text{adv}}=\mathcal{L_{\text{cls}}}(f(\tilde{\mathbf{E}}))$.
@@ 5 Experiments
Models. We conduct experiments on three MLLM familiies with different parameter size: Qwen2.5-vl-3b qwen25vl , LLaVA-1.5-7B & LLaVA-1.5-13B llava15 and InstructBLIP-Vicuna-7B instructblip . These models typically integrate a pre-trained vision encoder (e.g., CLIP ViT-L/14 clip ) with a language model backbone (e.g., Vicuna vicuna ), connected via a modality projection module. Following the original LLaVA and Qwen-VL implementation, we freeze the vision encoder and train the multimodal projector and the language model. In our adaptation of InstructBLIP, we adopt an alternative tuning strategy: we freeze both the vision encoder and the Q-Former, and fine-tune the language model and the linear projection layer. We provide further discussion in  Appendix B.
@@
Datasets. We evaluate models on benchmarks covering three task types: (i) Image-heavy tasks: Mini-ImageNet miniimagenet and Caltech-101 caltech101 , used in both training and evaluation, originally designed for image classification; (ii) Text-heavy tasks: OpenBookQA openbookqa and MMLU mmlu , consisting purely of textual question answering data; (iii) VQA tasks: For training, we use LLaVA-Instruct-dataset llava as the instruction-tuning dataset for related models. For InstructBLIP, we additionally include TextCaps textcaps , since these are the only publicly accessible VQA datasets known to be used in instruction-tuning for related models. For Qwen2.5-VL, whose instruction-tuning data is in-house and not publicly available, we use LLaVA-Instruct-dataset as a standardized alternative. For evaluation, we adopt three multiple-choice VQA benchmarks: ScienceQA-IMG scienceqa , MM-Bench-EN mmbench , and Seed-Bench-IMG seedbench . For ScienceQA and Seed-Bench, we only include examples with image context. For MM-Bench, we use the English version. All datasets are converted into a unified multiple-choice VQA format, enabling consistent modeling and evaluation across tasks and models. We report the accuracy of all multiple choice tasks and quantify the causal effect with the prediction change rate $\delta_{\text{cp}}$. All models are fine-tuned for 1 epoch with a fixed batch size $N_{\text{batch}}$. All the hyperparameters are listed in the  Appendix D.
@@
Model Settings Mini-ImageNet Caltech-101 OpenBookQA MMLU ScienceQA MM-Bench Seed-Bench
@@
As shown in Table 1, our method outperforms both VLMClassifier and I-MoF across different base MLLMs, demonstrating stronger robustness to modality interference and improved cross-modality competency. While I-MoF enhances visual grounding by integrating multiple visual features, it still suffers from modality interference: e.g. LLaVA-1.5-13B + I-MoF achieves 93.9% on original Mini-ImageNet but drops to 70.1% under perturbation (↓23.8%), indicating reliance on spurious textual cues. In contrast, our method maintains perturbed performance at 98.4% (↓0.1%). On the other hand, VLMClassifier, adopts vision-only fine-tuning, which leads to two critical limitations: vulnerability to cross-modal interference and degradation on VQA tasks, as LLaVA-1.5-7b + VLMClassifier-1 only reaches 35.8%/36.2% on MM-Bench/SeedBench, notably lower than both base LLaVA (72.1%/64.5%) and our method (73.7%/68.4%). These results highlight that vision-centric strategies, without addressing modality alignment, are insufficient for robust multimodal understanding. In text-heavy tasks such as OpenBookQA and MMLU, our method also achieves superior perturbed performance—82.1% vs. 64.5% on OpenBookQA, and 55.8% vs. 39.1% on MMLU (LLaVA-1.5-13B)—highlighting that addressing modality interference directly, rather than merely improving representations, is key to robust multimodal reasoning. Overall, unlike prior methods that often trade off between unimodal and multimodal performance, our method consistently improves both, achieving Pareto-optimality.
@@
Model Method ScienceQA-IMG MM-Bench-EN Seed-Bench-IMG VQA Overall
@@
To evaluate the effectiveness of each component in our framework, we conduct a comprehensive ablation study across multiple models and scales. We compare the pretrained models with the following strategies: FFT with $D^{\text{VQA}}$ (standard finetuning on VQA data), FFT with $D^{\text{AUG}}$ (supervised finetuning on mixed multi-task datasets with heuristic perturbations), FFT+KL/JS (adding consistency regularization on KL or JS divergence), FFT+RG (injecting random Gaussian noise instead of adversarial perturbation), FFT+PGD (FFT with heuristic perturbations and adversarial perturbations), and Ours (combining both perturbation-based data augmentation and consistency regularization).  Table 2 presents overall VQA performance across different models and training strategies, while, Table 3 evaluates model robustness under unimodal settings by comparing accuracy on the original inputs versus the worst-performing perturbation cases. Together, these results show the effectiveness of our method in improving both general VQA accuracy and robustness under modality interference. Table 4 reports results on unimodal tasks with all the perturbations.  Table 5 is full version of Table 3. Across all model families (Qwen2.5-VL, InstructBLIP, and LLaVA-1.5) and model sizes (3B/7B/13B), our method consistently achieves best overall performance, improving accuracy on both unimodal and multimodal benchmarks. For instance, it boosts overall VQA accuracy (e.g., +4.7% on LLaVA-1.5-7B, +14.9% on InstructBLIP-7B), but also enhances robustness to modality interference–improve the performance under perturbations by over 50% on image-heavy tasks(e.g. 17.5% $\rightarrow$ 98.3% on InstructBLIP-7B on Caltech101).  Figure 4 are radar chart visualizations. ( Table 4,  Table 5,  Figure 4 in  Appendix C)

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated impressive capabilities across tasks, yet they often exhibit difficulty in distinguishing task-relevant from irrelevant signals—particularly in tasks like Visual Question Answering (VQA)—which can lead to susceptibility to misleading or spurious inputs. We refer to this broader limitation as the Cross-Modality Competency Problem—the model’s inability to fairly evaluate all modalities. This vulnerability becomes more evident in modality-specific tasks—such as image classification or pure text question answering—where models are expected to rely solely on one modality. In such tasks, spurious information from irrelevant modalities often lead to significant performance degradation. We refer to this failure as Modality Interference, which serves as a concrete and measurable instance of the cross-modality competency problem, and we further design a perturbation-based causal diagnostic experiment to verify and quantify this problem. To mitigate modality interference, we propose a novel framework to finetune MLLMs, including perturbation-based data augmentations with both heuristic perturbations and adversarial perturbations via Projected Gradient Descent (PGD), and a consistency regularization strategy applying on model outputs with original and perturbed inputs. Experiments on multiple benchmark datasets (image-heavy, text-heavy and VQA tasks) and multiple model families with different scales demonstrate significant improvements in robustness and cross-modality competency, indicating our method’s effectiveness in boosting unimodal reasoning ability while enhancing performance on multimodal tasks. Our code and models can be found at: https://github.com/luisrui/Modality-Interference-in-MLLMs.
@@ 1 Introduction
Recent studies have attributed this phenomenon to a variety of symptoms arising during the vision-language alignment process, such as catastrophic forgetting imagewikiqa ; mof ; cogvlm ; vila , knowledge conflict wang2024mdpo ; vlmconflict , and spurious correlations more ; spurlens . Catastrophic forgetting has been identified as a key factor in visual degradation imagewikiqa ; mof , where multimodal tuning of MLLM overrides its pretrained visual features. Cross-modal knowledge conflict wang2024mdpo ; vlmconflict impairs pure-text reasoning, as models often produce inconsistent outputs when visual inputs are introduced, reflecting misaligned visual and textual parametric memories. Additionally, studies on spurious correlations more ; spurlens show that MLLMs tend to rely on superficial cross-modal cues rather than task-relevant grounding. While these symptoms shed light on MLLMs’ limitations, most works treat these issues in isolation. For instance, architectural issues such as shallow cross-modal fusion have been widely discussed: lightweight projectors in models like LLaVA llava fail to fully align vision and language representations, resulting in unstable modality reliance mof ; vlmconflict ; EECA . Others attribute performance bottlenecks to data limitations—insufficient modality-specific supervision leads to impaired visual decoding imagewikiqa and diminished language understanding vila . Inspired by these observations, our insight is to unify these challenges under a broader perspective: the model’s inability to identify and rely on the modality that contributes most relevant information to the task. We argue that the fundamental limitation lies in MLLMs’ lack of cross-modality competency Competency —the ability to fairly evaluate and integrate information across modalities. Current MLLMs lack mechanisms to support this competency during inference, making them vulnerable to misleading cross-modal signals—a failure mode we refer to as Modality Interference.
@@
To systematically diagnose and mitigate modality interference, we introduce a two-stage methodology grounded in causal analysis. First, we design a perturbation-based evaluation experiment inspired by causal intervention principles more to diagnose the extent of modality interference across tasks and model scales. Second, we propose a robust fine-tuning framework to mitigate modality interference. Specifically, in our evaluation analysis, we first focus on modality-heavy settings using a multiple-choice question answering format, where the model selects an answer from a fixed set of options based on both image and text input. We then include image-heavy tasks (e.g., image classification), text-heavy tasks (e.g., pure-text QA), and balanced multimodal tasks (e.g., VQA), allowing us to examine how models behave under different modality-reliance scenarios. To further induce modality interference, we introduce heuristic perturbations: In image-heavy tasks, we perturb the text input by prepending either (i) unrelated scientific facts or (ii) misleading descriptions that falsely associate an incorrect option with the image content. In text-heavy tasks, where the default visual input is random noise, we perturb the visual input with (i) semantically meaningful real images, (ii) full black canvas, or (iii) full white canvas. These perturbations are designed to either introduce spurious cues or reinforce irrelevant modality signals. We evaluate the resulting changes in model predictions to assess the robustness of modality selectivity. While the perturbation-based evaluations offer empirical insights, we further frame our analysis through a causal intervention framework and in which we model modality interference through a causal graph abstraction. Building on this framework, we evaluate a range of pretrained MLLMs across different architectures and scales with results shown in Figure˜1. In image-heavy tasks, unrelated textual facts moderately reduce performance, while misleading descriptions cause severe degradation—revealing the model’s vulnerability to spurious textual cues. In text-heavy tasks, canvas inputs have little effect, but unrelated real images significantly hurt performance, indicating improper fusion of irrelevant visual signals into reasoning.
@@
The empirical results from Figure˜1 confirm the presence of modality interference and reveal the limitations of current MLLMs in lack of cross-modality competency. To mitigate modality interference, we propose a perturbation-based fine-tuning framework for MLLMs. Specifically, to alleviate data insufficiency, we apply a perturbation-based data augmentations strategy, where we construct a diverse mixture of modality-specific, perturbation-augmented samples and original VQA samples. The perturbations includes both heuristic perturbations (e.g. add unrelated facts in image-heavy prompts) and adversarial training time perturbations, which exposes the model to worst-case alignment disruptions, serving as a strong variant. To further improve robustness, we introduce a consistency regularization strategy(e.g., via Jensen–Shannon divergence), which enforces output stability between original and perturbed samples. In summary, the main contributions of this paper are threefold. First, we introduce the notion of the Cross-Modality Competency Problem to describe how multimodal models may struggle to balance different modalities, and analyze modality interference as one concrete instance of such challenges in MLLMs. Second, we design a perturbation-based causal evaluation experiment that systematically quantifies modality reliance and reveals models’ susceptibility to modality interference. Third, we propose a fine-tuning strategy that combines supervised augmentation with both heuristic and adversarial perturbations and consistency regularization to mitigate modality interference. Extensive experiments across multiple MLLM families llava15 ; instructblip ; qwen25vl and diverse benchmarks miniimagenet ; caltech101 ; openbookqa ; mmlu ; scienceqa ; mmbench ; seedbench demonstrate the superiority of our method.
@@ 3 Causal Analysis on Modality Interference
Competency problems describe scenarios where models rely on spurious correlations between isolated input features and output labels to make predictions, instead of leveraging meaningful interactions among multiple features Competency . We extend this concept to the multimodal setting by treating entire modalities (e.g., image $X_{I}$ or text $X_{T}$) as structured feature sources. We define the Cross-Modality Competency as an ability for MLLM which no single modality should be sufficient to determine the correct prediction, and models are expected to integrate information across modalities to make robust decisions. A competent model should be able to fairly evaluate and integrate all modalities, identifying which ones carry task-relevant signals while ignoring misleading or irrelevant ones. For instance, in a pure-text question answering task, the model receives both a question and an image, as is standard in MLLM input formats. However, the image is not required to answer the question. If the model relies on spurious visual cues—such as objects or scenes that frequently co-occur with certain answers—it violates the task’s competency condition by grounding predictions in irrelevant modality signals. This manifests as Modality Interference, where the presence of an irrelevant but misleading modality disrupts the model’s reasoning.
@@
To systematically measure cross-modality competency, we propose a perturbation-based evaluation framework. The core idea is to inject controlled noise into the irrelevant modality and assess the model’s robustness to such perturbations. Specifically, for image-heavy tasks, we perturb the text input by: (1)Prepending unrelated scientific facts; (2) Prepending misleading descriptions that falsely link incorrect options to the image content. For text-heavy tasks, we perturb the visual input by: (1) Attaching a real but irrelevant image; (2) Substituting with a full black or full white canvas image. Models with strong modality selectivity should maintain high prediction consistency when irrelevant modality signals are perturbed. We select multiple image-heavy and text-heavy tasks for evaluation. Each task is framed as a multiple-choice classification problem, requiring the model to choose the correct option based on image and text modalities as input, with perturbations applied as described above. Details in Section˜5 and Section˜C.1.
@@
We formalize modality interference through a causal intervention perspective with a causal graph, as shown in Figure˜2, where visual inputs ($X_{I}$) and textual inputs ($X_{T}$) are processed into their respective representations ($Z_{I}$, $Z_{T}$) before being fused to produce the final prediction ($A$). To study the model’s reliance on different modalities, we introduce perturbations directly at the input level, serving as causal interventions causalIntervention on $X_{I}$ and $X_{T}$. Under ideal cross-modality competency, the model’s prediction should primarily depend on the task-relevant pathway (e.g., $X_{I}\rightarrow Z_{I}\rightarrow A$ in image-heavy tasks, $X_{T}\rightarrow Z_{T}\rightarrow A$ in text-heavy tasks). Causal interventions at the input level allow us to diagnose whether the model improperly fuses irrelevant signals into its decision process. We use $x_{I}\textquoteright$ to denote an intervention on image $X_{I}$ and use $x_{T}\textquoteright$ as an intervention on text $X_{T}$. Following Pearl’s causal framework causalIntervention ; intermediaryVariable , we quantify the impact of modality perturbations on model predictions by formalizing causal effects in our multimodal setting. Specifically, we define the pre-intervention prediction distribution as $P(A|X_{I},X_{T})$, and the post-intervention prediction distribution after applying a perturbation on $X_{I}$ or $X_{T}$ as $P\textquoteright(A|\text{do}(X_{I}=x_{I}\textquoteright)$ or $\text{do}(X_{T}=x_{T}\textquoteright))$. The do-operation represents an intervention to specific modality, and the causal effect (CE) of an intervention is evaluated via a distance metric $\delta$ comparing $P$ and $P\textquoteright$ as $\text{CE}=\delta(P,P^{\prime})$. We assess the causal effect via prediction changes using $\delta_{\text{cp}}(P,P^{\prime}):=\mathbb{I}\left(a\neq a^{\prime}\right)$ in which $a=\arg\max_{x}P(x)$ is the predicted answer before intervention, $a\textquoteright=\arg\max_{x}P\textquoteright(x)$ is the predicted answer after intervention and $\mathbb{I}(\cdot)$ is the indicator function that outputs $1$ if $a\neq a\textquoteright$ and 0 otherwise. Thus, $\delta_{\text{cp}}$ captures whether the model’s final decision $A$ changes under perturbations to the input modality. In all interventions, a high value of $\delta_{\text{cp}}$ indicates the model’s susceptibility to modality interference, revealing spurious reliance on irrelevant modality.
@@ 4 Methods
To mitigate modality interference and enhance cross-modality competency, we propose a unified perturbation-aware training framework that introduces interventions at both the input level (on $X_{I}$ and $X_{T}$) and the representation level (on $Z_{I}$ and $Z_{T}$) with consistency regularization. Overall pipeline is displayed in  Figure˜3.
@@ 4.1 Perturbation-based Data Augmentation
We sample $N_{\text{img}}$ and $N_{\text{text}}$ examples from $\mathcal{D}^{\text{img}}$ and $\mathcal{D}^{\text{text}}$ to construct $\mathcal{B}^{\text{img}}_{\text{orig}}$ and $\mathcal{B}^{\text{text}}_{\text{orig}}$ respectively, and the remaining $N_{\text{vqa}}$ examples are VQA samples from $\mathcal{D}^{\text{vqa}}$ to construct $\mathcal{B}^{\text{vqa}}$ . With dynamically generated perturbed variants for each image-heavy and text-heavy sample, the final training batch is:
@@
where $\alpha$ is the step size and $\Pi$ projects the noise back into the $\ell_{\infty}$ ball with radius $\epsilon$, we initialize $\delta$ with Gaussian noise $\mathcal{N}(0,\epsilon^{2})$. After obtaining $\tilde{\mathbf{E}}$, we compute the adversarial loss as $)\mathcal{L}_{\text{adv}}=\mathcal{L_{\text{cls}}}(f(\tilde{\mathbf{E}}$)).
@@ 5 Experiments
Models. We conduct experiments on three MLLM familiies with different parameter size: Qwen2.5-vl-3b qwen25vl , LLaVA-1.5-7B & LLaVA-1.5-13B llava15 and InstructBLIP-Vicuna-7B instructblip . These models typically integrate a pre-trained vision encoder (e.g., CLIP ViT-L/14 clip ) with a language model backbone (e.g., Vicuna vicuna ), connected via a modality projection module. Following the original LLaVA and Qwen-VL implementation, we freeze the vision encoder and train the multimodal projector and the language model. In our adaptation of InstructBLIP, we adopt an alternative tuning strategy: we freeze both the vision encoder and the Q-Former, and fine-tune the language model and the linear projection layer. We provide further discussion in  Appendix˜B.
@@
Datasets. We evaluate models on benchmarks covering three task types: (i) Image-heavy tasks: Mini-ImageNet miniimagenet and Caltech-101 caltech101 , used in both training and evaluation, originally designed for image classification; (ii) Text-heavy tasks: OpenBookQA openbookqa and MMLU mmlu , consisting purely of textual question answering data; (iii) VQA tasks: For training, we use LLaVA-Instruct-dataset llava as the instruction-tuning dataset for related models. For InstructBLIP, we additionally include TextCaps textcaps , since these are the only publicly accessible VQA datasets known to be used in instruction-tuning for related models. For Qwen2.5-VL, whose instruction-tuning data is in-house and not publicly available, we use LLaVA-Instruct-dataset as a standardized alternative. For evaluation, we adopt three multiple-choice VQA benchmarks: ScienceQA-IMG scienceqa , MM-Bench-EN mmbench , and Seed-Bench-IMG seedbench . For ScienceQA and Seed-Bench, we only include examples with image context. For MM-Bench, we use the English version. All datasets are converted into a unified multiple-choice VQA format, enabling consistent modeling and evaluation across tasks and models. We report the accuracy of all multiple choice tasks and quantify the causal effect with the prediction change rate $\delta_{\text{cp}}$. All models are fine-tuned for 1 epoch with a fixed batch size $N_{\text{batch}}$. All the hyperparameters are listed in the  Appendix˜D.
@@
Model Settings Mini-ImageNet Caltech-101 OpenBookQA MMLU ScienceQA $M$-Bench Seed-Bench
@@
As shown in Table˜1, our method outperforms both VLMClassifier and I-MoF across different base MLLMs, demonstrating stronger robustness to modality interference and improved cross-modality competency. While I-MoF enhances visual grounding by integrating multiple visual features, it still suffers from modality interference: e.g. LLaVA-1.5-13B + I-MoF achieves 93.9% on original Mini-ImageNet but drops to 70.1% under perturbation (↓23.8%), indicating reliance on spurious textual cues. In contrast, our method maintains perturbed performance at 98.4% (↓0.1%). On the other hand, VLMClassifier, adopts vision-only fine-tuning, which leads to two critical limitations: vulnerability to cross-modal interference and degradation on VQA tasks, as LLaVA-1.5-7b + VLMClassifier-1 only reaches 35.8%/36.2% on MM-Bench/SeedBench, notably lower than both base LLaVA (72.1%/64.5%) and our method (73.7%/68.4%). These results highlight that vision-centric strategies, without addressing modality alignment, are insufficient for robust multimodal understanding. In text-heavy tasks such as OpenBookQA and MMLU, our method also achieves superior perturbed performance—82.1% vs. 64.5% on OpenBookQA, and 55.8% vs. 39.1% on MMLU (LLaVA-1.5-13B)—highlighting that addressing modality interference directly, rather than merely improving representations, is key to robust multimodal reasoning. Overall, unlike prior methods that often trade off between unimodal and multimodal performance, our method consistently improves both, achieving Pareto-optimality.
@@
Model Method ScienceQA-IMG $M$-Bench-EN Seed-Bench-IMG VQA Overall
@@
To evaluate the effectiveness of each component in our framework, we conduct a comprehensive ablation study across multiple models and scales. We compare the pretrained models with the following strategies: FFT with $D^{\text{VQA}}$ (standard finetuning on VQA data), FFT with $D^{\text{AUG}}$ (supervised finetuning on mixed multi-task datasets with heuristic perturbations), FFT+KL/JS (adding consistency regularization on KL or JS divergence), FFT+RG (injecting random Gaussian noise instead of adversarial perturbation), FFT+PGD (FFT with heuristic perturbations and adversarial perturbations), and Ours (combining both perturbation-based data augmentation and consistency regularization).  Table˜2 presents overall VQA performance across different models and training strategies, while, Table˜3 evaluates model robustness under unimodal settings by comparing accuracy on the original inputs versus the worst-performing perturbation cases. Together, these results show the effectiveness of our method in improving both general VQA accuracy and robustness under modality interference. Table˜4 reports results on unimodal tasks with all the perturbations.  Table˜5 is full version of Table˜3. Across all model families (Qwen2.5-VL, InstructBLIP, and LLaVA-1.5) and model sizes (3B/7B/13B), our method consistently achieves best overall performance, improving accuracy on both unimodal and multimodal benchmarks. For instance, it boosts overall VQA accuracy (e.g., +4.7% on LLaVA-1.5-7B, +14.9% on InstructBLIP-7B), but also enhances robustness to modality interference–improve the performance under perturbations by over 50% on image-heavy tasks(e.g. 17.5% $\rightarrow$ 98.3% on InstructBLIP-7B on Caltech101).  Figure˜4 are radar chart visualizations. ( Table˜4,  Table˜5,  Figure˜4 in  Appendix˜C)